given a standard deck of 52 cards. what is the probability that drawing one card: you will draw a ten card that is black?

Answer by fractalier(6550)

P(ten) = 4/52 = 1/13
P(black) = 26/52 = 1/2
P(ten AND black) = 1/13 * 1/2 = 1/26